The week of July 11

Last week we reported on a incident with the HET tracker that put us in a hard skew condition and shut us down. We had recovered from that skew condition but the software issues that led to the skew had continued. This week I am happy to report that we have not only installed additional skew sensors but we have found the source of the software issues. The software teams in Austin and West Texas were able to perform a series of experiments which allowed them to isolate the problem as coming from the ethernet card inside the dSpace box. This is a very specialized ethernet card but thankfully we were able to swap out a different card for it and found that all of our TCS communications problems, hexapod faults, and loading trajectory errors have gone away. We will look into getting a spare card and send the card in to see what kind of fault occurred in the card.

With the improved tracker performance we were able to start our early science operations. We took our first science targets on the night of the 16th. Unfortunately, we are now headed towards full moon which is not the ideal time to do science targets with LRS2. However, it does give us a chance to work out any inefficiencies in our PI interface, night operations interfaces and queue tools so that we can be fully ready once we get some dark time a few days after full moon.

Leave a Reply

Your email address will not be published. Required fields are marked *